OpenStudy (anonymous):

integral of (1 + cos x)^2

OpenStudy (anonymous):

i guess you have to square it and take \[\int(1+2\cos(x)+\cos^2(x))dx\]

myininaya (myininaya):

\[x+2\sin(x)+\int\limits_{}^{}\frac{1}{2}(1+\cos(2x)) dx +C\]

OpenStudy (anonymous):

first to are rather simple and for the second one use \[\cos^2(x)=\frac{1}{2}\cos(2x)+\frac{1}{2}\]

OpenStudy (anonymous):

a standard trick for \[\cos^2(x)\] so standard that i might be worth memorizing

OpenStudy (anonymous):

i bet myininaya has it memorized for sure right?

myininaya (myininaya):

or you could use cos(x+x)=cos(x)cos(x)-sin(x)sin(x) and sin^2(x)+cos^2(x)=1 to derive it

OpenStudy (anonymous):

sure if you had an infinite amount of time

myininaya (myininaya):

lol it doesn't take long

OpenStudy (anonymous):

but since it never changes might as well learn it at least until calc class is over. kind of like knowing that \[7\times 8=56\] or that \[\int\ln(x)dx=x\ln(x)-x\]